OpenStudy (anonymous):

IS ANYOME GOOD AT FINDING DERIATIVES FOR LN?! Please I need help as soon as possible! :(

OpenStudy (alekos):

post it up

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

y=ln(1-x/(x+2)^5)

OpenStudy (alekos):

\[y=\ln [1 -\frac{ x }{(x+2)^{5}}]\] Is that it?

OpenStudy (anonymous):

No sorry my computer is messed up, the 1-x is the numerator, and the denominator is as you out it. Ln is correct too

OpenStudy (anonymous):

Put*

OpenStudy (anonymous):

@alekos

OpenStudy (irishboy123):

\(\large y=ln \frac{1-x}{(x+2)^5} \) assuming that you know that \(\large (log \ u(x))' = \frac{1}{u} u'\) then look also at \(\large ln \frac{a}{b} = ln \ a - ln \ b\) and similar properties of logs; and maybe split this up a bit before you start...

OpenStudy (alekos):

yeah, thanks Irish Boy. I reckon do the properties of logs.... Ln(1-x) - Ln(x+2)^5 = Ln(1-x) - 5Ln(x+2) the rest is easy because the derivative of Lnx is 1/x

OpenStudy (irishboy123):

@alekos look good